# Thermal one-loop self-energy correction for hydrogen-like systems: relativistic approach

### v1

@@ I Introduction
Given the modern development of spectroscopic experiments and the significantly increased accuracy in determining transition frequency standards over recent decades, increasingly smaller effects in terms of their absolute value are beginning to play a key role. To achieve such progress, the evolution of quantum theory and the corresponding methods and approaches was required. The main tool in the study of the spectroscopic properties of atomic systems is the theory of quantum electrodynamics (QED) [Greiner, AB], which allows for an unprecedented comparison of theoretical and experimental results. A clear example for such analysis can be the hydrogen atom, the accuracy of measurement of the transition frequency in which reaches an order of $10^{-15}$ relative value, see, e.g., [CODATA-2021]. Attention can also be drawn to the use of various hydrogen-like ions and other one-electron systems, with which such physical quantities as the g-factor of a bound electron and consequently the electron mass [Sturm2014], the proton-to-electron mass ratio, etc., [Karr_2025] have been determined with a percentage per trillion level of accuracy.
@@
Covering a wide range of tasks, the comparative analysis of theory with experimental measurement results typically reduces to calculating QED corrections of the next order of smallness (within various perturbation theories, different smallness parameters may be used). At present, it can be stated that for light atomic systems, theoretical research already deals with corrections of the order of $\alpha^{7}$ ($\alpha$ is the fine structure constant), see, for example [Korobov2014, PPY-2017, CODATA-2021]. Aiming for precise determination of fundamental physical constants, verification of fundamental interactions, or investigation of physical symmetries violations, this kind of analysis occasionally reveals differences that are then used to construct theoretical hypotheses beyond the framework of the Standard Model [PPY-2017, PhysRevA.98.022501, Kozlov-RevModPhys, Blaum_2021].
@@
However, the primary standard for achieving the highest accuracy in spectroscopic measurements is atomic clocks, as well as various candidates for this role, see [Kozlov-RevModPhys] and the references therein. Atomic clocks, which allow for highly precise measurements, have become a sensitive tool for detecting changes in the gravitational field [Bothwell_2022], testing global symmetry [Sanner2019], or imposing constraints on dark matter [DM-atcl]. Experimental precision at the level of $10^{-18}$ and beyond has necessitated theoretical studies of effects that cause even the slightest shifts in atomic energy levels, and also disrupt the stability of the operational transition in atomic clocks.
@@
It is precisely these effects that reflect the influence of the thermal environment on atomic characteristics. Blackbody radiation (BBR) was the first such influence to be observed experimentally [GC], with estimates for the resulting frequency shifts and transition rates provided therein. A more rigorous quantum-electrodynamical treatment followed in [farley]. The theory and approaches presented in [farley] for calculating the thermal Stark effect and the corresponding transition probabilities induced by BBR are fundamental even today. Such effects have drawn considerable attention in recent decades, effectively becoming a distinct area of theoretical research. Forming one of the dominant contributions to the overall experimental error in determining physical quantities, efforts have been directed towards both the most precise calculations (see, e.g., the works by M. Safronova and co-authors [safronova_b_al_in, AtCl-Sr, Saf-Nat] as well as more resent [Beloy_2014, PhysRevA.110.043108, Beloy_2025]), and the search for ways to mitigate the influence of BBR [Yudin, Zuhrianda, PhysRevA.100.023417].
@@
Despite various possibilities for suppressing thermal shifts of atomic energy levels, the main theoretical task remains their accurate calculation. Even to offset the dominant contribution requires meticulous control of temperature fluctuations and, as a consequence, determining the resulting uncertainties. As noted above, the primary approach for such calculations is based on quantum mechanical theory [farley]. This approach, extended to many-body perturbation theory, has been successfully applied to estimate the dc-Stark shift and its dynamic corrections induced by blackbody radiation [Saf-Nat, Sahoo-AtCl]. However, around the same time as the development of QM theory of BBR influence on atomic characteristics, the theory of quantum electrodynamics at finite temperature (TQED) was also evolving [Dol, Don, DHR]. This theory encompasses a much broader range of problems related to the thermal environment, and allows for a precise description of phenomena even outside of a perturbative regime with respect to temperature [Soto, Soto2].
@@
In this work, focusing on atomic bound states, we consider temperature ranges relevant to laboratory conditions—specifically, those well below the ionization threshold of the atomic electron. Previously, it was shown that to obtain the well-known quantum mechanical (QM) results for the ac-Stark shift and the depopulation rates of atomic levels, it is sufficient to take into account the thermal one-loop self-energy (SE) correction for the bound electron [PRA_2015, S_2020]. Then the dominant contribution (Stark shift) arises within the nonrelativistic limit for the operators written in the dipole approximation (longwave approximation). Then the dominant contribution (dynamic Stark shift) arises in the non-relativistic limit for the operator written in the dipole approximation (long-wavelength approximation).
@@
Demonstrating excellent agreement with the QM approach, the TQED result for the SE correction, nonetheless, is not limited to just the Stark shift. A comprehensive analysis of this approach was presented very recently in work [Lopez_2025]. Specifically, it was found that in addition to the dipole dynamic polarizability of the atom, this correction also contains higher multipole polarizabilities, including magnetic, quadrupole, etc. (similar to [Porsev]), as well as relativistic corrections arising from the wave functions and the expansion of the self-energy operator. Thus, a fully relativistic approach to calculating the thermal one-loop self-energy correction for a bound electron appears more advantageous for obtaining an accurate value of the atomic energy level shift. Additionally, it’s worth noting that the TQED approach readily allows for the consideration of even more subtle effects, whether they relate to hyperpolarizabilities or corrections to physical quantities such as the g-factor or hyperfine structure in various atomic systems, see, e.g., [china-BBR, ZGSg_2022, ZGS_2022, ZKS_2023, ZS_2023].
@@
In this study, we provide a comprehensive analysis of the thermal one-loop self-energy correction of a bound electron within a fully relativistic approach. Discarding the imaginary part (corresponding to the BBR-induced level width [PRA_2015]), we focus on precise calculations of the real part of such a correction for the hydrogen atom. Using the partial wave decomposition method, we analyze the behavior of this thermal correction as the nuclear charge $Z$ changes. Calculations are presented for various states of the hydrogen atom (up to a principal quantum number of $n=12$ and an orbital quantum number of $l\leq 3$) at room temperature, and for various temperatures for states that more closely correspond to the most accurately measured transition frequencies [CODATA-2021]. Throughout this article the relativistic units ($\hbar=c=m=1$, where $m$ is the electron mass) are used.
@@ II.1 QED description of thermal shifts and line broadening
Referring to the QED description of the atomic level energy shift caused by BBR, we rely on the main conclusions drawn in [Lopez_2025]. According to this theory, the most significant thermal contributions can be obtained by considering the one-loop self-energy correction of the bound electron, in which the ordinary Feynman photon propagator is replaced by the corresponding thermal part [Dol, Don, DHR]. Specifically, it was shown in [PRA_2015] that the real part of such correction provides the dominant contribution to the atomic level energy shift (Stark shift), while the imaginary part corresponds to the line broadening caused by blackbody radiation. The Feynman diagram showing the thermal one-loop SE correction is illustrated in Fig. 1.
@@
In Eq. (II.1), $S(x_{1},x_{2})$ denotes the electron propagator, which admits an eigenmode expansion in terms of the bound-state spectrum, as detailed in, e.g., [Greiner]:
@@
In turn, the photon propagator $D_{\mu\nu}(x_{1},x_{2})$ can be used in the form [PRA_2015, S_2020]:
@@
Omitting intermediate calculations for brevity (see, for example, [PRA_2015]), the following expression for the thermal SE correction can be obtained:
@@
Here ${\boldsymbol{\alpha}}_{i}$ ($i=1,2$) are the Dirac alpha matrices acting in the $i$-th coordinate space. Taking into account the imaginary additive in the energy denominator, using the Sokhotski–Plemelj theorem, it is easy to show that the Stark shift arises from the real part, $\Re\Delta\varepsilon_{a}^{\beta}$, and the line broadening caused by BBR from the imaginary part, $\Im\Delta\varepsilon_{a}^{\beta}$, of Eq. (6), see [PRA_2015, S_2020, Lopez_2025].
@@ II.2 Partial-wave expansion
To perform the rigorous relativistic calculations (in contrast to the non-relativistic limit and its extensions as in [Lopez_2025]), we use the partial wave decomposition, see [abr], for the operator $I(\omega)$:
@@ III Technical details, numerical results and discussion
Below we evaluate the real part of the thermal one-loop self-energy correction numerically for excited states of the hydrogen atom and the ground state of various hydrogen-like ions. The Dirac equation is solved using the finite basis set method. The basis set is obtained from the expansion of the wave function in terms of B-splines within the framework of the dual-kinetic-balance (DKB) approach [DKB]. This method excludes the so-called spurious states and establishes the correct asymptotics of the wave functions in the non-relativistic limit.
@@
First, we present the results for the calculations of the energy shift of the ground state of hydrogen-like ions with the arbitrary nuclear charge $Z$. The real part of the thermal SE correction is calculated according to Eq. (10). The obtained results for a temperature of $300$ K are listed in Table 1 and illustrated by the graph in Fig. 2, which shows the dependence of the energy shift on the nuclear charge $Z$. The parametric estimate for the real part of Eq. (10) can be achieved in the standard way: $r\sim 1/(m\alpha Z)$, $\varepsilon\sim m(\alpha Z)^{2}$ in relativistic units. Then, in the dipole approximation, $I(|\omega|)\sim\bm{r}_{1}\bm{r}_{2}$, see [Lopez_2025], combined with the static limit, one can find: $\Delta\varepsilon_{a}^{\beta}\sim(k_{\rm B}T)^{4}/m^{3}\alpha^{3}Z^{4}$, i.e., $\Delta\varepsilon_{a}^{\beta}\sim 1/Z^{4}$.
Table 1: The real part of the one-loop self-energy correction for the $1s_{1/2}$ state of hydrogen-like ions in Hz at a temperature of $T=300$ K. For $Z<10$, a spherical nuclear model is used, and for $Z\geq 10$, a Fermi nuclear model is used. Results from Ref. [farley] are marked with an asterisk , from Ref. [PRA_2015] with a cross , and from Ref. [reex] with two asterisks ∗∗. The estimate of the numerical calculation error is indicated in parentheses, and the order of the calculated shift is presented separately for convenience.
@@
Despite the fact that the numerical calculation worsens with an increasing nuclear charge, the deviation from the asymptotic behavior of $1/Z^{4}$ can be explained by relativistic corrections to the wave functions and the self-energy operator, as well as higher multipole moments [Lopez_2025]. The estimate $\Delta\varepsilon_{a}^{\beta}$ for large $Z$ confirms that the thermal shift becomes negligible in highly relativistic hydrogen-like systems. Based on this, we focus mainly on light hydrogen-like systems.
The Table 2 indicates the real part of the self-energy correction for the excited states of hydrogen $\Re\Delta\varepsilon_{a}^{\beta}$. Within the framework of our fully relativistic method, the values for the thermal dynamic effect Eqs. (6), (7) align with good accuracy to the Stark shift obtained using a purely quantum mechanical approach [farley]. The most significant discrepancy (around $8\%$ for the $2s$ state) arises from the fine structure of the levels, as well as from the Lamb shift and the numerical calculation method, which did not allow for the determination of as many significant figures as indicated in Table 2; these factors were not accounted for in [farley, PRA_2015, jetp2022]. In turn, comparison with [reex] shows excellent agreement of the results. We should emphasize that the accuracy of our results can be confirmed by calculations of the thermal Zeeman shift (see below) without additional analytical derivations [Lopez_2025] and carefully accounts for the fine structure within a fully relativistic approach (in contrast to [jetp2022]).
Table 2: Real part of the self-energy correction for the excited states of hydrogen $\Re\Delta\varepsilon_{a}^{\beta}$ in Hz. The temperature of $300$ K is considered. Results with included Lamb shift are highlighted in italics. Numbers from Ref. [farley] are marked with an asterisk , from Ref. [PRA_2015] with a dagger , and from Ref. [jetp2022] with two asterisks ∗∗.

### v2

@@ I Introduction
Given the modern development of spectroscopic experiments and the significantly increased accuracy in determining transition frequency standards over recent decades, increasingly smaller effects in terms of their absolute value are beginning to play a key role. To achieve such progress, the evolution of quantum theory and the corresponding methods and approaches was required. The main tool in the study of the spectroscopic properties of atomic systems is the theory of quantum electrodynamics (QED) [17, 2], which allows for an unprecedented comparison of theoretical and experimental results. A clear example for such analysis can be the hydrogen atom, the accuracy of measurement of the transition frequency in which reaches an order of $10^{-15}$ relative value, see, e.g., [46]. Attention can also be drawn to the use of various hydrogen-like ions and other one-electron systems, with which such physical quantities as the g-factor of a bound electron and consequently the electron mass [44], the proton-to-electron mass ratio, etc., [22] have been determined with a percentage per trillion level of accuracy.
@@
Covering a wide range of tasks, the comparative analysis of theory with experimental measurement results typically reduces to calculating QED corrections of the next order of smallness (within various perturbation theories, different smallness parameters may be used). At present, it can be stated that for light atomic systems, theoretical research already deals with corrections of the order of $\alpha^{7}$ ($\alpha$ is the fine structure constant), see, for example [23, 29, 46]. Aiming for precise determination of fundamental physical constants, verification of fundamental interactions, or investigation of physical symmetries violations, this kind of analysis occasionally reveals differences that are then used to construct theoretical hypotheses beyond the framework of the Standard Model [29, 11, 24, 6].
@@
However, the primary standard for achieving the highest accuracy in spectroscopic measurements is atomic clocks, as well as various candidates for this role, see [24] and the references therein. Atomic clocks, which allow for highly precise measurements, have become a sensitive tool for detecting changes in the gravitational field [7], testing global symmetry [35], or imposing constraints on dark matter [47]. Experimental precision at the level of $10^{-18}$ and beyond has necessitated theoretical studies of effects that cause even the slightest shifts in atomic energy levels, and also disrupt the stability of the operational transition in atomic clocks.
@@
It is precisely these effects that reflect the influence of the thermal environment on atomic characteristics. Blackbody radiation (BBR) was the first such influence to be observed experimentally [15], with estimates for the resulting frequency shifts and transition rates provided therein. A more rigorous quantum-electrodynamical treatment followed in [14]. The theory and approaches presented in [14] for calculating the thermal Stark effect and the corresponding transition probabilities induced by BBR are fundamental even today. Such effects have drawn considerable attention in recent decades, effectively becoming a distinct area of theoretical research. Forming one of the dominant contributions to the overall experimental error in determining physical quantities, efforts have been directed towards both the most precise calculations (see, e.g., the works by M. Safronova and co-authors [33, 28, 32] as well as more resent [4, 45, 5]), and the search for ways to mitigate the influence of BBR [49, 56, 27].
@@
Despite various possibilities for suppressing thermal shifts of atomic energy levels, the main theoretical task remains their accurate calculation. Even to offset the dominant contribution requires meticulous control of temperature fluctuations and, as a consequence, determining the resulting uncertainties. As noted above, the primary approach for such calculations is based on quantum mechanical theory [14]. This approach, extended to many-body perturbation theory, has been successfully applied to estimate the dc-Stark shift and its dynamic corrections induced by blackbody radiation [32, 34]. However, around the same time as the development of QM theory of BBR influence on atomic characteristics, the theory of quantum electrodynamics at finite temperature (TQED) was also evolving [8, 10, 9]. This theory encompasses a much broader range of problems related to the thermal environment, and allows for a precise description of phenomena even outside of a perturbative regime with respect to temperature [12, 13].
@@
In this work, focusing on atomic bound states, we consider temperature ranges relevant to laboratory conditions—specifically, those well below the ionization threshold of the atomic electron. Previously, it was shown that to obtain the well-known quantum mechanical (QM) results for the ac-Stark shift and the depopulation rates of atomic levels, it is sufficient to take into account the thermal one-loop self-energy (SE) correction for the bound electron [42, 43]. Then the dominant contribution (Stark shift) arises within the nonrelativistic limit for the operators written in the dipole approximation (longwave approximation). Then the dominant contribution (dynamic Stark shift) arises in the non-relativistic limit for the operator written in the dipole approximation (long-wavelength approximation).
@@
Demonstrating excellent agreement with the QM approach, the TQED result for the SE correction, nonetheless, is not limited to just the Stark shift. A comprehensive analysis of this approach was presented very recently in work [26]. Specifically, it was found that in addition to the dipole dynamic polarizability of the atom, this correction also contains higher multipole polarizabilities, including magnetic, quadrupole, etc. (similar to [31]), as well as relativistic corrections arising from the wave functions and the expansion of the self-energy operator. Thus, a fully relativistic approach to calculating the thermal one-loop self-energy correction for a bound electron appears more advantageous for obtaining an accurate value of the atomic energy level shift. Additionally, it’s worth noting that the TQED approach readily allows for the consideration of even more subtle effects, whether they relate to hyperpolarizabilities or corrections to physical quantities such as the g-factor or hyperfine structure in various atomic systems, see, e.g., [55, 51, 52, 53, 54].
@@
In this study, we provide a comprehensive analysis of the thermal one-loop self-energy correction of a bound electron within a fully relativistic approach. Discarding the imaginary part (corresponding to the BBR-induced level width [42]), we focus on precise calculations of the real part of such a correction for the hydrogen atom. Using the partial wave decomposition method, we analyze the behavior of this thermal correction as the nuclear charge $Z$ changes. Calculations are presented for various states of the hydrogen atom (up to a principal quantum number of $n=12$ and an orbital quantum number of $l\leq 3$) at room temperature, and for various temperatures for states that more closely correspond to the most accurately measured transition frequencies [46]. Throughout this article the relativistic units ($\hbar=c=m=1$, where $m$ is the electron mass) are used.
@@ II.1 QED description of thermal shifts and line broadening
Referring to the QED description of the atomic level energy shift caused by BBR, we rely on the main conclusions drawn in [26]. According to this theory, the most significant thermal contributions can be obtained by considering the one-loop self-energy correction of the bound electron, in which the ordinary Feynman photon propagator is replaced by the corresponding thermal part [8, 10, 9]. Specifically, it was shown in [42] that the real part of such correction provides the dominant contribution to the atomic level energy shift (Stark shift), while the imaginary part corresponds to the line broadening caused by blackbody radiation. The Feynman diagram showing the thermal one-loop SE correction is illustrated in Fig. 1.
@@
In Eq. (II.1), $S(x_{1},x_{2})$ denotes the electron propagator, which admits an eigenmode expansion in terms of the bound-state spectrum, as detailed in, e.g., [17]:
@@
In turn, the photon propagator $D_{\mu\nu}(x_{1},x_{2})$ can be used in the form [42, 43]:
@@
Omitting intermediate calculations for brevity (see, for example, [42]), the following expression for the thermal SE correction can be obtained:
@@
Here ${\boldsymbol{\alpha}}_{i}$ ($i=1,2$) are the Dirac alpha matrices acting in the $i$-th coordinate space. Taking into account the imaginary additive in the energy denominator, using the Sokhotski–Plemelj theorem, it is easy to show that the Stark shift arises from the real part, $\Re\Delta\varepsilon_{a}^{\beta}$, and the line broadening caused by BBR from the imaginary part, $\Im\Delta\varepsilon_{a}^{\beta}$, of Eq. (6), see [42, 43, 26].
@@ II.2 Partial-wave expansion
To perform the rigorous relativistic calculations (in contrast to the non-relativistic limit and its extensions as in [26]), we use the partial wave decomposition, see [1], for the operator $I(\omega)$:
@@ III Technical details, numerical results and discussion
Below we evaluate the real part of the thermal one-loop self-energy correction numerically for excited states of the hydrogen atom and the ground state of various hydrogen-like ions. The Dirac equation is solved using the finite basis set method. The basis set is obtained from the expansion of the wave function in terms of B-splines within the framework of the dual-kinetic-balance (DKB) approach [41]. This method excludes the so-called spurious states and establishes the correct asymptotics of the wave functions in the non-relativistic limit.
@@
First, we present the results for the calculations of the energy shift of the ground state of hydrogen-like ions with the arbitrary nuclear charge $Z$. The real part of the thermal SE correction is calculated according to Eq. (10). The obtained results for a temperature of $300$ K are listed in Table 1 and illustrated by the graph in Fig. 2, which shows the dependence of the energy shift on the nuclear charge $Z$. The parametric estimate for the real part of Eq. (10) can be achieved in the standard way: $r\sim 1/(m\alpha Z)$, $\varepsilon\sim m(\alpha Z)^{2}$ in relativistic units. Then, in the dipole approximation, $I(|\omega|)\sim\bm{r}_{1}\bm{r}_{2}$, see [26], combined with the static limit, one can find: $\Delta\varepsilon_{a}^{\beta}\sim(k_{\rm B}T)^{4}/m^{3}\alpha^{3}Z^{4}$, i.e., $\Delta\varepsilon_{a}^{\beta}\sim 1/Z^{4}$.
Table 1: The real part of the one-loop self-energy correction for the $1s_{1/2}$ state of hydrogen-like ions in Hz at a temperature of $T=300$ K. For $Z<10$, a spherical nuclear model is used, and for $Z\geq 10$, a Fermi nuclear model is used. Results from Ref. [14] are marked with an asterisk , from Ref. [42] with a cross , and from Ref. [21] with two asterisks ∗∗. The estimate of the numerical calculation error is indicated in parentheses, and the order of the calculated shift is presented separately for convenience.
@@
[…]
black The numerical calculation error given in parentheses in Table I is determined as follows. First, the partial-wave decomposition in Eq. (8) constitutes a rapidly convergent series owing to the small value of the parameter $\omega$, which is defined by the Planck distribution (temperature). Extrapolation of the partial-wave series is not required and this effect can therefore be disregarded when estimating the calculation error. We then performed calculations with different sets of basis functions to determine the convergence of the result. The stated mean value was chosen considering a sufficiently fast computation time while ensuring a convergent result with the number of basis functions set to $140$ B-splines, with convergence already observed at $100$ B‑splines. Another parameter is the box size (was finally set to $600$ atomic units), which allows one to control the number of recovered ’real’ atomic bound states. Recall that the DKB method replaces the atomic spectrum with an effective, purely discrete set of states. Because the continuous spectrum mixes with the discrete energy levels in this case, only the lowest few states correspond well to the ’real’ ones. Their number depends on the box size and can be determined by comparing the Dirac energy and the pseudolevel energy. Thus, by varying these two parameters, we have, in our view, achieved optimal convergence of the obtained values with the corresponding stated numerical error. The same consideration regarding the pseudo-spectrum results in a larger numerical error for highly excited states (see below). We also note that, in practice, all numerical calculations (those presented above and subsequently) were carried out to machine precision.
@@
Despite the fact that the numerical calculation worsens with an increasing nuclear charge, the deviation from the asymptotic behavior of $1/Z^{4}$ can be explained by relativistic corrections to the wave functions and the self-energy operator, as well as higher multipole moments [26]. The estimate $\Delta\varepsilon_{a}^{\beta}$ for large $Z$ confirms that the thermal shift becomes negligible in highly relativistic hydrogen-like systems. Based on this, we focus mainly on light hydrogen-like systems.
The Table 2 indicates the real part of the self-energy correction for the excited states of hydrogen $\Re\Delta\varepsilon_{a}^{\beta}$. Within the framework of our fully relativistic method, the values for the thermal dynamic effect Eqs. (6), (7) align with good accuracy to the Stark shift obtained using a purely quantum mechanical approach [14]. The most significant discrepancy (around $8\%$ for the $2s$ state) arises from the fine structure of the levels, as well as from the Lamb shift and the numerical calculation method, which did not allow for the determination of as many significant figures as indicated in Table 2; these factors were not accounted for in [14, 42, 50]. In turn, comparison with [21] shows excellent agreement of the results. We should emphasize that the accuracy of our results can be confirmed by calculations of the thermal Zeeman shift (see below) without additional analytical derivations [26] and carefully accounts for the fine structure within a fully relativistic approach (in contrast to [50]).
Table 2: Real part of the self-energy correction for the excited states of hydrogen $\Re\Delta\varepsilon_{a}^{\beta}$ in Hz. The temperature of $300$ K is considered. Results with included Lamb shift are highlighted in italics. Numbers from Ref. [14] are marked with an asterisk , from Ref. [42] with a dagger , and from Ref. [50] with two asterisks ∗∗.